OpenStudy (anonymous):

If f(x) = 5x2 – 8x + 13 and g(x) = 3x – 8, what is f(x) + g(x)?

OpenStudy (anonymous):

5x^2

OpenStudy (precal):

add the functions together

OpenStudy (anonymous):

8^x

OpenStudy (agent47):

Just add them

mathslover (mathslover):

5x^2-8x+13+3x-8 =f(x)+g(x)

mathslover (mathslover):

since f(x) = 5x^2-8x+13 and g(x) = 3x-8 hence f(x)+g(x) = add them

OpenStudy (anonymous):

5 x^2 - 5x +5

OpenStudy (anonymous):

Thankyou!!

OpenStudy (anonymous):

my pleasure :)